Show that x=1-(b/a) satisfies the equation a(a+b)x 2 -b(a-b)x=(a-b) 2 ,and find the other value of x.


Show that  x=1-(b/a) satisfies the equation a(a+b)x2-b(a-b)x=(a-b)2,and find the other value of x.


Grade:9

2 Answers

Chetan Mandayam Nayakar
12 years ago

x=(a-b)/a, a(a+b)(a-b)2/a2 -(b(a-b)(a-b)/a)=((a-b)/a)=((a-b)/a)(a2-b2+b2-ab)=a(a-b)(a-b)/a =(a-b)2

                other value = (a-b)2/(a(a+b))*(a/(a-b))= (a-b)/(a+b)

sanatan sharma
12 years ago

x=(a-b)/a,

a(a+b)(a-b)2/a2 -(b(a-b)(a-b)/a)=(a-b)2

=((a-b)/a)=(a-b)2

=((a-b)/a)(a2-b2+b2-ab)=(a-b)2

=a(a-b)(a-b)/a =(a-b)2

α+β=[-(-b(a-b))]/[a(a+b)]

(a-b)/a +β = [b(a-b)]/[a(a+b)]

β = (a-b)[{b/a(a+b)} - 1/a]

therefore β = (b-a)/(b+a)